# Learned Indexes with Distribution Smoothing via Virtual Points

### v1

@@ Abstract
Recent research on learned indexes has created a new perspective for indexes as models that map keys to their respective storage locations. These learned indexes are created to approximate the cumulative distribution function of the key set, where using only a single model may have limited accuracy. To overcome this limitation, a typical method is to use multiple models, arranged in a hierarchical manner, where the query performance depends on two aspects: (i) traversal time to find the correct model and (ii) search time to find the key in the selected model. Such a method may cause some key space regions that are difficult to model to be placed at deeper levels in the hierarchy. To address this issue, we propose an alternative method that modifies the key space as opposed to any structural or model modifications. This is achieved through making the key set more learnable (i.e., smoothing the distribution) by inserting virtual points. Further, we develop an algorithm named CSV to integrate our virtual point insertion method into existing learned indexes, reducing both their traversal and search time. We implement CSV on state-of-the-art learned indexes and evaluate them on real-world datasets. The extensive experimental results show significant query performance improvement for the keys in deeper levels of the index structures at a low storage cost.
Index Terms:
Learned indexes, Distribution smoothing, Index optimisation.
I Introduction
@@
Learned indexes [1] have reported strong query performance and are attracting much attention from both the academia and industry in recent years. The core idea of learned indexes is that an index structure can be seen as a mapping function from search keys to storage locations of data records. The mapping function (a.k.a. indexing function) is learned and approximated by machine learning algorithms (models). To enable the learning, some storage ordering needs to be established. Typically, an ascending order based on the search keys is used, such that the mapping function is effectively the cumulative distribution function (CDF) of the search keys.
@@
Different learned indexes have been proposed [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], with a common theme to design indexing functions and structures that enable better approximation of the CDF, since approximation errors translate to mapping errors and hence extra search costs to recover from the errors. However, this approach means the use of either complex indexing functions (e.g., splines [18, 19]) or piece-wise functions with many segments [3, 12, 20], both of which could lead to sub-optimal query efficiency. This is illustrated by Fig. 1 with LIPP [4] – one of the latest learned indexes that places “difficult to learn” keys in deeper levels of the index. As the figure shows, keys indexed in deeper levels (i.e., higher levels in the figure) reported higher query times on average, on all four datasets consistently.
@@
Figure 1: Query time at each level of the LIPP index for four real datasets, each with 200 million keys.
@@
Our core idea is to add virtual points to “smooth” the CDF of a dataset to be indexed. Take Fig. 2(a) as an example, where each black dot represents a data point (i.e., its search key). Approximating the CDF of the dataset with a linear indexing function can result in a large approximation error (and hence high search costs at query time) for keys $k_{1}$ and $k_{2}$. We sum up the squared prediction error of every point:
$\mathcal{L}_{f}(K)=\sum_{i=1}^{n}\big{(}f(k_{i})-rank(k_{i})\big{)}^{2},$ (1)
where $K$ denotes the set of keys and $n$ is its size, $k_{i}\in K$ is a key, $rank(k_{i})$ is its rank, and $f(\cdot)$ denotes the indexing function, respectively. We refer to $\mathcal{L}_{f}(K)$ as sum of squared errors (SSE). In this case, $\mathcal{L}_{f}(K)=8.33$ – a large value of $\mathcal{L}_{f}(K)$ suggests worse prediction accuracy using $f$ for search key mapping and hence higher query times.
As Fig. 2(b) shows, we add a set of virtual points $V=\{k_{v1},k_{v2},\ldots,k_{v5}\}$ represented by the red hollow dots. Here, we assume a smoothing budget of $50.5n=$ 5, i.e., 5 virtual points are allowed. Now the original data points are spread out, and the CDF of the (original and virtual) points is closer to a straight line. We refit the points with a new index function $f^{\prime}$, with the SSE $\mathcal{L}_{f^{\prime}}(K)$ being reduced to 2.05 (SSE with the new virtual points, $\mathcal{L}_{f^{\prime}}(K\cup V)$ = 2.29).
@@
Figure 2: Indexing data points (keys) with CDF smoothing.
We show that, given a smoothing budget that constraints the additional space costs, finding the optimal placement of the virtual points to minimise the SSE $\mathcal{L}_{f^{\prime}}(K)$ is NP-hard. We then consider approximation solutions for two generic scenarios: (1) smoothing the CDF for index learning with a single indexing function and (2) smoothing the CDF for a hierarchy of indexing functions, which is a common structure of existing learned indexes. We focus on linear index functions for their high efficiency, although the idea of CDF smoothing extends to more complex (e.g., quadratic) functions naturally.
We propose an algorithm named CDF smoothing via virtual points (CSV) to smooth the CDF for optimising hierarchical learned indexes, with the aim to reduce the overall height of the structures as well as the prediction errors of each indexing function, and hence the query costs. This is performed by collecting sub-trees of the hierarchical structure and smoothing the CDF of the keys in them by inserting virtual points. As a result, the keys could now be placed into a single node due to the higher learnability, provided it surpasses a cost model threshold value. Here, the cost model is used to balance the reduction in query traversal time and the potential increase in the leaf-node search time due to the increase of keys in a node.
It is important to note that our aim is not to propose yet another learned index but rather a technique that can be integrated with existing or emerging hierarchical learned indexes to optimise their query efficiency with controllable extra space. To show the applicability of our CSV algorithm, we integrate it with three recent learned indexes ALEX [2], LIPP [4], and SALI [14], which are the state-of-the-art (SOTA). Experimental results show that, powered by CSV, these learned indexes have a consistent improvement in the query efficiency, with limited overheads in the storage space.
@@
  • We propose a key space transformation technique using CDF smoothing via inserting virtual points to enhance the index learnability.
@@
  • We propose an efficient algorithm named CSV to integrate the CDF smoothing technique with hierarchical learned indexes to improve the query performance of such indexes, with a controllable space overhead.
@@
  • We integrate CSV with three learned indexes ALEX, LIPP, and SALI, and we conduct experiments with four real datasets. The experimental results show that the learned indexes powered by CSV manage to promote up to 60% of the keys in lower levels to upper levels, resulting in up to 34% improvement of their query time.
II Related Work
@@
We first review learned indexes in general. Then, we discuss the gapped array technique used to accommodate insertions and their relevance to our work. Afterwards, we focus on studies addressing complex distributions, which share a similar goal with us. We also cover a technical called the poisoning attacks, which motivates our CDF smoothing technique.
II-A Learned Indexes
@@
Learned indexes are a trending topic in the database community [2, 3, 4, 5, 9, 1, 18, 15, 16]. Their key insight is to treat indexes as functions that map a search key to the storage position of the corresponding data object, which can be learned using machine learning models. A common approach is to layout the data objects by ascending order of their search keys, such that the indexing functions are effectively (approximations of) CDFs of the search keys.
@@
To index large datasets, multiple indexing functions are used, typically organized in a hierarchical structure like a B-tree. The lookup performance of such a structure is then dominated by two steps: (1) the traversal time to find the leaf-node (every node corresponds to an indexing function) indexing the search key, and (2) the search time within the selected leaf-node (leaf-node search time hereafter to distinguish from the traversal time) to locate the target data object, as the indexing functions have prediction errors and may not produce the exact storage position of the search target [21, 22].
@@
It is a challenge to balance the query costs from these two steps mentioned above. While a deeper structure with more indexing functions may fit the data distribution better and have lower leaf-node search times, it may also have higher traversal times and larger index size [23, 24]. Some studies impose a maximum error bound on the indexing functions to reduce the leaf-node search times [20, 3], also at the cost of more indexing functions. Another approach is to use more complex indexing functions (as opposed to linear ones) [1, 13, 18], e.g., splines, which could better fit the CDFs. The issue with this approach is the higher function inference time, and hence higher query and insertion times [21].
@@
These studies design structures and indexing functions to better fit the data distribution. We address the challenge from an alternate perspective, i.e., we adjust the data distribution such that it is easier to be fit by the indexing functions.
II-B Learned Indexes with Gapped Arrays
Several learned indexes [2, 4, 15] leave gaps in their storage structure (i.e., gapped arrays). While their purpose is to accommodate data insertions, a side effect is changing the data distribution, which is what we do. A core difference to note is that, they have not considered minimising the indexes’ model prediction errors when adding gaps, while we do.
II-C Learned Indexes Addressing Complex Data Distributions
@@
To better index CDFs of complex data distributions, there are two common approaches. One is to use more complex functions such as splines and piece-wise linear regression models [3, 18]. The other is to use better data partitioning strategies for easier CDF learning over each partition, such as by CARMI [25] and EWALI [26]. Another study, LER [27], uses logarithmic error-based loss functions (instead of the more commonly used least squared error-based) to improve the learning of index models that better fit the CDF.
@@
A latest development, SALI [14], identifies the most frequently accessed nodes via probability models given a query workload. The corresponding sub-trees are flattened by using a segmentation approach, similar to the PGM index [3], to reduce the traversal time within these sub-trees. However, this leads to an additional search step for query processing, as we need to find the correct node from the flattened structure.
@@
A couple of studies [23, 28] transform the input key set into a more uniform distribution to improve the CDF learnability. The NFL index [23] transforms the key distribution using a numerical normalizing flow that transforms a latent distribution to a new distribution via generative models. The distribution transformation introduces overheads, while queries also need to be transformed to use the index. Further, the transformation may increase the tail conflict degree for certain distributions, making it unsuitable in those instances. The gap insertion [28] technique inserts gaps between the keys (i.e., storage positions of the corresponding data objects) to straighten the CDF of the keys, thereby improving its learnability. However, this is performed by manipulating the rank of each key, and as a result, multiple keys could be given the same position. An extra array is used to house such conflicting keys, which in turn introduces search overheads to locate the correct key at query time. Further, this method leads to a heavy storage space increase of up to 87%. Neither NFL nor gap insertion considers future insertions at construction time. They are less resilient against data updates.
II-D Poisoning Cumulative Distribution Functions
@@
Our idea of adjusting data distribution to fit the indexing functions is rooted from data poisoning – a process of manipulating the training data to change the results from a predictive model [29]. Data poisoning has been introduced into learned indexes to poison the indexing functions and negatively impact their capability to approximate the CDFs [29]. The main goal of this process is to identify new points to include into the original key set that would cause the maximum increase to the loss function value (i.e., the SSE).
@@
Motivated by the poisoning technique, we propose a technique that smooths the data distribution by adding virtual points (i.e., new keys), to obtain CDFs that are easier to be approximated by indexing functions (models), hence leading to a structure with higher query efficiency. Since the models are built with virtual points that can be used to host data insertions, a side benefit of our structure is that it is more resilient against data insertions. Table I highlights the key difference between our technique (CSV), NFL and gap insertion.
TABLE I: Comparison with Existing Works
CSV NFL [23] Gap insertion [28]
@@ III Preliminaries
Suppose that the index keys have been partitioned and indexed by a set $\mathcal{F}$ of $m$ indexing functions. Each indexing function $f_{i}\in\mathcal{F}$ has some prediction error for a key $k\in K_{i}$ indexed by it. Here, $K_{i}\subset K$ refers to the subset of keys indexed by $f_{i}$. The prediction error refers to the squared difference between the predicted index position $f_{i}(k)$ and the rank of $k$ in $K$, i.e., $rank(k)$.
Let $\mathcal{L}_{\mathcal{F}}$ be the total sum of squared errors of all indexing functions in $\mathcal{F}$:
$\mathcal{L}_{\mathcal{F}}(K)=\sum_{i=1}^{m}\sum_{k\in K_{i}}\big{(}f_{i}(k)-rank(k)\big{)}^{2}.$ (2)
The sum of squared errors (SSE) is chosen as it is one of the most commonly used metric to represent loss in the existing studies of learned indexes.
@@
Our aim is to insert values (virtual points) into $K$ while keeping it sorted, i.e., to smooth the CDF of $K$, such that the total SSE is minimised.
@@
A naive optimal smoothing scheme is to insert as many virtual points as needed such that every point $k\in K_{i}$ lies at the $f_{i}(k)$-th position (i.e., $rank(k)=f_{i}(k)$) in the list (assuming unique integer keys). This way, the total prediction error becomes zero after smoothing. In reality, this smoothing scheme is not always feasible, due to the non-uniqueness of the keys in $K$ and the potentially high space cost.
@@
We consider a “smoothing budget” $\lambda$, i.e., the number of virtual points allowed to be inserted, such that total SSE is minimised given the constraint of $\lambda$.
Definition 1
@@
[Learned index smoothing] Given a list of index keys $K$ sorted in ascending order and partitioned into $m$ segments, each of which is indexed by an indexing function $f_{i}\in\mathcal{F}$, the learned index smoothing problem aims to insert a set $V$ ($|V|\leq\lambda$) of virtual points into $K$ while keeping $K$ in order, such that the total SSE is minimised.
@@
We consider linear indexing functions as they are used in most existing learned indexes. We assume $\lambda=\alpha\cdot n$ where the smoothing threshold $\alpha$ is in $(0,1)$, to retain a linear space overhead. To simplify the discussion, we use integer index keys, while our techniques also apply to real number index keys when they can be scaled up to become integers.
@@
NP-hardness analysis. Solving the exact CDF smoothing problem is NP-hard because it can be reduced from the min-Knapsack problem, which is a known NP-hard problem. The min-Knapsack problem considers a given set of items $S$ and a target weight $t$. Each item $s\in S$ is associated with a cost $c_{s}$ and a weight $w_{s}$. The objective is determining the subset $A\subseteq S$ that minimises the total cost of the items in $A$ while the total weight of the items is equal to or greater than $t$.
@@
Our learned index smoothing problem considers a key set $K=\{k_{1},k_{2},\ldots,k_{n}\}$ of size $n$. We aim to find a subset $V$ (virtual points, $k_{vi}$) of at most size $\lambda$ from a candidate set $C$ that would minimise the loss function value $\mathcal{L}$. Naively, the set $C$ can be formed by considering $\lambda$ virtual point candidates between every two adjacent keys in $K$, i.e., $|C|\leq\lambda\cdot(n-1)$.
To reduce the NP-hard min-Knapsack to our problem, we set the weight of every item to 1. Further, the target weight value $t$ should be set to 0, as it is possible to have no virtual points to reduce the loss value. To enforce the smoothing budget, we add an upper limit $\lambda$ to the summation of the weights, which would not change the NP-hardness of this min-Knapsack problem. Now the set $S$ can be mapped to set $C$, and choosing the best virtual point subset $V$ can be mapped to finding the subset $A$ that minimises the loss function in our problem and the objective function in the min-Knapsack problem, respectively. This transformation can be performed in polynomial time, and when our problem is solved, the min-Knapsack problem is solved. As such, our problem is NP-hard.
@@
Due to the difficulties in finding an exact optimal solution for the learned index smoothing problem, next, we consider two reduced versions of the problem and propose highly effective heuristic solutions: (1) smoothing the CDF for the subset of keys $K_{i}$ indexed by an indexing function $f_{i}$ (Section IV); and (2) smoothing the CDFs for all $m$ subsets $K_{i}$ when they are indexed under a hierarchical learned index (Section V).
IV CDF Smoothing for a Single Linear Model

### v2

@@ Abstract.
Recent research on learned indexes has created a new perspective for indexes as models that map keys to their respective storage locations. These learned indexes are created to approximate the cumulative distribution function of the key set, where using only a single model may have limited accuracy. To overcome this limitation, a typical method is to use multiple models, arranged in a hierarchical manner, where the query performance depends on two aspects: (i) traversal time to find the correct model and (ii) search time to find the key in the selected model. Such a method may cause some key space regions that are difficult to model to be placed at deeper levels in the hierarchy. To address this issue, we propose an alternative method that modifies the key space as opposed to any structural or model modifications. This is achieved through making the key set more learnable (i.e., smoothing the distribution) by inserting virtual points. Furthermore, we develop an algorithm named CSV to integrate our virtual point insertion method into existing learned indexes, reducing both their traversal and search time. We implement CSV on state-of-the-art learned indexes and evaluate them on real-world datasets. Extensive experimental results show significant query performance improvement for the keys in deeper levels of the index structures at a low storage cost.
Learned indexes, Distribution smoothing, Index optimisation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data access methods
1. Introduction
@@
Learned indexes (Kraska et al., 2018) have reported strong query performance and are attracting much attention from both the academia and industry in recent years. The core idea of learned indexes is that an index structure can be seen as a mapping function $f(\cdot)$ from a search key $k_{i}$ to the storage location (i.e., the rank $rank(k_{i})$) of the corresponding data record: $rank(k_{i})\approx f(k_{i})$. The mapping function (a.k.a. indexing function) is learned and approximated by machine learning algorithms (models). To enable the learning, a storage ordering needs to be established. Typically, an ascending order based on the search keys is used, such that the mapping function is effectively the cumulative distribution function (CDF) of the search keys.
@@
Different learned indexes have been proposed (Ding et al., 2020a; Ferragina and Vinciguerra, 2020; Wu et al., 2021; Zhang et al., 2021; Nathan et al., 2020; Li et al., 2020; Ding et al., 2020b, 2022; Pai et al., 2024; Sheng et al., 2023; Li et al., 2021b; Tang et al., 2020; Ge et al., 2023b; Lu et al., 2021; Lan et al., 2023; Wang et al., 2024), with a common theme to design indexing functions and structures that enable better approximation of the CDF, since approximation errors translate to mapping errors (i.e., difference between $rank(k_{i})$ and $f(k_{i})$) and hence extra search costs to examine data records at around $f(k_{i})$ and recover from the errors. However, such approaches mean the use of either complex indexing functions (e.g., splines (Kipf et al., 2020; Mishra and Singhal, 2021)) or piece-wise functions with many segments (Ferragina and Vinciguerra, 2020; Li et al., 2021b; Galakatos et al., 2019), both of which could lead to sub-optimal query efficiency. This is illustrated by Fig. 1 with LIPP (Wu et al., 2021) – one of the latest learned indexes. The index has a hierarchical structure built in a top-down manner. When an index model (i.e., a node in the index) cannot achieve an overall low mapping error for all keys assigned to it for indexing, sub-index models are created recursively as a hierarchy, to accommodate for the “more difficult to learn” keys. As Fig. 1 shows, keys indexed in deeper levels (higher levels in the figure) reported higher query times on average, on all four datasets.
@@
Figure 1. Query time at each level of the LIPP index for four real datasets, each with 200 million keys.
@@
Our core idea is to add virtual points to “smooth” the CDF of a dataset. Take Fig. 2(a) as an example, where each black dot represents a data point (i.e., its search key). Approximating the CDF of the dataset with a linear function can result in a large approximation error (and hence high search costs at query time) for keys $k_{1}$ and $k_{2}$. We sum up the squared prediction error of every point:
(1) $\mathcal{L}_{f}(K)=\sum_{i=1}^{n}\big{(}f(k_{i})-rank(k_{i})\big{)}^{2},$
where $K$ denotes the set of keys and $n$ is its size, $k_{i}\in K$ is a key, $rank(k_{i})$ is its rank, and $f(\cdot)$ is the indexing function. We refer to $\mathcal{L}_{f}(K)$ as the loss function, for which we use the sum of squared errors (SSE). In this case, $\mathcal{L}_{f}(K)=8.33$ – a large value of $\mathcal{L}_{f}(K)$ suggests worse prediction accuracy using $f$ for search key mapping and hence higher query times.
As Fig. 2(b) shows, we can add virtual points $V=\{k_{v1},k_{v2},\ldots,k_{v5}\}$ represented by the red hollow dots. Here, we assume a smoothing budget of $50.5n=$ 5, i.e., 5 virtual points are allowed. Now the original data points are spread out, and the CDF of the (original and virtual) points is closer to a straight line. We refit the points with a new indexing function $f^{\prime}$, with the loss function value $\mathcal{L}_{f^{\prime}}(K)$ being reduced to 2.04 (and $\mathcal{L}_{f^{\prime}}(K\cup V)$ = 2.29).
@@
Figure 2. Indexing data points (keys) with CDF smoothing.
We show that, given a smoothing budget that constrains the additional space costs, finding the optimal placement of the virtual points to minimise the loss function $\mathcal{L}_{f^{\prime}}(K)$ is NP-hard. We then propose approximation solutions for two generic scenarios: (1) smoothing the CDF for index learning with a single indexing function and (2) smoothing the CDF for a hierarchy of indexing functions, which is a common structure for existing learned indexes. We focus on linear indexing functions for their efficiency, although CDF smoothing can naturally extend to more complex (e.g., quadratic) functions.
We propose an algorithm named CDF smoothing via virtual points (CSV) to smooth the CDF for optimising hierarchical learned indexes, with the aim to reduce the overall height of the structures as well as the prediction errors of each indexing function, and hence the query costs. This is performed by collecting sub-trees of the hierarchical structure and smoothing the CDF of the keys in them. As a result, the keys could now be placed into a single node due to the higher learnability, provided it surpasses a cost model threshold value. Here, the cost model is used to balance the reduction in index traversal time and the potential increase in the leaf-node search time due to the increase of keys in a node.
It is important to note that our aim is not to propose yet another learned index but rather a technique that can be integrated with existing or emerging hierarchical learned indexes to optimise their query efficiency with controllable extra space. To show the applicability of our CSV algorithm, we integrate it with three recent learned indexes ALEX (Ding et al., 2020a), LIPP (Wu et al., 2021), and SALI (Ge et al., 2023b), which are the state-of-the-art (SOTA).
@@
(1) We propose a key space transformation technique using CDF smoothing via inserting virtual points to enhance index learnability.
@@
(2) We propose an efficient algorithm named CSV to integrate the CDF smoothing technique with hierarchical learned indexes to improve the query performance, with a controllable space overhead.
@@
(3) We integrate CSV with three learned indexes ALEX, LIPP, and SALI, and we conduct experiments with four real datasets. The experimental results show that the learned indexes powered by CSV manage to promote up to 60% of the keys in lower levels to upper levels, resulting in up to 34% improvement of their query time, with less than 15% increase to the storage space overhead.
2. Related Work
@@
We first review learned indexes in general. Then, we focus on studies addressing complex distributions, which share a similar goal with us. We also cover a technique called the poisoning attacks, which motivates our CDF smoothing technique.
2.1. Learned Indexes
@@
Learned indexes are a trending topic in the database community (Ding et al., 2020a; Ferragina and Vinciguerra, 2020; Wu et al., 2021; Zhang et al., 2021; Ding et al., 2022; Kraska et al., 2018; Kipf et al., 2020; Lu et al., 2021; Yang et al., 2023; Lan et al., 2023). Their key idea is to treat indexes as functions that map a search key to the storage position of the corresponding data object, which can be learned with machine learning models. A common approach is to lay out the data objects by ascending order of their search keys, such that the indexing functions are effectively (approximations of) CDFs of the keys.
@@
To index large datasets, multiple indexing functions are used, typically organized in a hierarchical structure like a B-tree. The lookup performance of such a structure is then dominated by two steps: (1) the traversal time to find the leaf-node (every node corresponds to an indexing function) indexing the search key, and (2) the search time within the selected leaf-node (leaf-node search time hereafter to distinguish from the traversal time) to locate the target data object, as the indexing functions have errors and may not produce the exact storage position of the search target (Sun et al., 2023; Wongkham et al., 2022).
@@
It is a challenge to balance the query costs from the two steps above. While a deeper structure with more indexing functions may fit the data distribution better and have lower leaf-node search times, it may also have higher traversal times and larger index sizes (Wu et al., 2022; Ge et al., 2023a). Some studies impose a maximum error bound on the indexing functions to reduce the leaf-node search times (Galakatos et al., 2019; Ferragina and Vinciguerra, 2020), also at the cost of more indexing functions. Another approach is to use more complex indexing functions (as opposed to linear ones) (Kraska et al., 2018; Tang et al., 2020; Kipf et al., 2020), e.g., splines, which could better fit the CDFs. The issue with this approach is the higher inference time for the function, and hence higher query and insertion times (Sun et al., 2023).
@@
These studies design structures and indexing functions to better fit the data distribution. We address the challenge from an alternate perspective, i.e., we adjust the data distribution such that it is easier to be fitted by the indexing functions.
2.2. Addressing Complex Data Distributions
@@
To better index CDFs of complex data distributions, there are two common approaches. One is to use more complex functions such as splines and piece-wise linear regression models (Ferragina and Vinciguerra, 2020; Kipf et al., 2020). The other is to use better data partitioning strategies for easier CDF learning over each partition, such as by CARMI (Zhang and Gao, 2022) and EWALI (Liu et al., 2023). Another study, LER (Eppert et al., 2021), uses logarithmic error-based loss functions (instead of the more commonly used least squared error-based) to improve the learning of index models that better fit the CDF.
@@
A latest development, SALI (Ge et al., 2023b), identifies the most frequently accessed nodes via probability models given a query workload. The corresponding sub-trees are flattened using a segmentation approach, similar to the PGM index (Ferragina and Vinciguerra, 2020), to reduce their traversal time. However, this leads to an additional search step for queries, as we need to find the correct node from the flattened structure.
@@
A couple of studies (Wu et al., 2022; Li et al., 2021a) transform the input key set into a more uniform distribution to improve the CDF learnability. The NFL index (Wu et al., 2022) transforms the key distribution using a numerical normalizing flow that transforms a latent distribution to a new distribution via generative models. The distribution transformation introduces overheads, while queries also need to be transformed to use the index. Further, the transformation may increase the tail conflict degree for certain distributions, making it unsuitable in those instances. The gap insertion (Li et al., 2021a) technique inserts gaps between the keys (i.e., storage positions of the corresponding data objects) to straighten the CDF of the keys, thereby improving its learnability. However, this is performed by manipulating the rank of each key, and as a result, multiple keys could be given the same position. An extra array is used to house such conflicting keys, which in turn introduces search overheads to locate the correct key. This method leads to a heavy storage space increase of up to 87%.
Several learned indexes (Ding et al., 2020a; Wu et al., 2021; Lu et al., 2021) leave gaps in their storage structure (i.e., gapped arrays). While their purpose is to accommodate data insertions, a side effect is changing the data distribution, which is what we do. A core difference to note is that, they do not consider minimizing the indexes’ model prediction errors when adding gaps, in contrast to our approach which does.
2.3. Poisoning CDFs
@@
Our idea of adjusting data distribution to fit the indexing functions is rooted from data poisoning – a process of manipulating the training data to change the results from a predictive model (Kornaropoulos et al., 2022). Data poisoning has been introduced into learned indexes to poison the indexing functions and negatively impact their capability to approximate the CDFs (Kornaropoulos et al., 2022). The main goal of this process is to identify new points to include into the original key set that would cause the maximum increase to the loss function value (i.e., the SSE).
@@
Motivated by the poisoning technique, we propose a technique that smooths the data distribution by adding virtual points, to obtain CDFs that are easier to be approximated by indexing functions (models), hence leading to a structure with higher query efficiency. Since the models are built with virtual points that can be used to host data insertions, a side benefit of our structure is that it is more resilient against data insertions. Table 1 highlights the key difference between our technique CSV, NFL, and gap insertion (GI).
Table 1. Comparison with Existing Works
CSV NFL (Wu et al., 2022) GI (Li et al., 2021a)
@@ 3. Preliminaries
Suppose that the index keys have been partitioned and indexed by a set $\mathcal{F}$ of $m$ indexing functions. Each indexing function $f_{i}\in\mathcal{F}$ has some prediction error (a.k.a “loss”) for a key $k\in K_{i}$ indexed by it. Here, $K_{i}\subset K$ refers to the subset of keys indexed by $f_{i}$. The loss refers to the squared difference between the predicted index position $f_{i}(k)$ and the rank of $k$ in $K$, i.e., $rank(k)$. The sum of squared errors (SSE) is one of the most commonly used metric to represent loss in the existing studies of learned indexes. Let $\mathcal{L}_{\mathcal{F}}$ be the total sum of squared errors of all indexing functions in $\mathcal{F}$:
(2) $\mathcal{L}_{\mathcal{F}}(K)=\sum_{i=1}^{m}\sum_{k\in K_{i}}\big{(}f_{i}(k)-rank(k)\big{)}^{2}.$
@@
Equation 2 is the loss function of our optimisation problem. We aim to insert values (virtual points) into $K$ while keeping it sorted, such that $\mathcal{L}_{\mathcal{F}}(K)$ is minimised, i.e., to smooth the CDF of $K$.
@@
A naive optimal smoothing scheme is to insert as many virtual points as needed such that every point $k\in K_{i}$ lies at the $f_{i}(k)$-th position (i.e., $rank(k)=f_{i}(k)$) in the list (assuming unique integer keys). This way, the loss becomes zero after smoothing. In reality, this smoothing scheme is infeasible, due to the non-uniqueness of the keys in $K$ and the potentially high space cost. For example, if the keys are 64-bit integers, it will take $2^{64}\times 8$ bytes $\approx 128$ exabyte (i.e., $128\times 10^{6}$ TB) to achieve such an index key layout.
@@
To balance between the space overhead and the smoothness of the CDF with inserted points, we consider a “smoothing budget” $\lambda$, i.e., the number of virtual points allowed to be inserted, such that the loss is minimised given the constraint of $\lambda$. We assume $\lambda=\alpha\cdot n$ where the smoothing threshold $\alpha$ is in $(0,1)$, to retain a linear space overhead. Formally, we aim to solve the following problem:
Definition 0.
@@
[Learned index smoothing] Given a list of index keys $K$ sorted in ascending order and partitioned into $m$ segments, each of which is indexed by an indexing function $f_{i}\in\mathcal{F}$, the learned index smoothing problem aims to insert a set $V$ ($|V|\leq\lambda$) of virtual points into $K$ while keeping $K$ in order, such that the loss as defined by Equation 2 is minimised.
@@
We consider linear indexing functions as they are used in most existing learned indexes. To simplify the discussion, we use integer index keys, while our techniques also apply to real number index keys when they can be scaled up to become integers.
NP-hardness analysis. Solving the exact CDF smoothing problem is NP-hard as it can be reduced from the Knapsack problem, which is a known NP-hard problem.
Lemma 3.2.
Learned index CDF smoothing is NP-hard.
Proof.
@@
We reduce from the Knapsack problem which is NP-hard. The Knapsack problem considers a set of items $S$. Each item $s\in S$ has a value $c_{s}$ and a weight $w_{s}$. The objective is to determine the subset $A\subseteq S$ that maximises the total value of the items in $A$ while the total weight of the items is less than a given limit $t$.
@@
CDF smoothing considers a key set $K$ of size $n$. We aim to find a subset $V$ (virtual points $k_{vi}$) of at most size $\lambda$ from a candidate set $C$ that would minimise the loss $\mathcal{L}$ (which implies the maximisation of loss reduction from the loss without CDF smoothing). Naively, the set $C$ can be formed by considering $\lambda$ virtual point candidates between every two adjacent keys in $K$, i.e., $|C|\leq\lambda\cdot(n-1)$.
To reduce the Knapsack problem to our CDF smoothing problem, the set $S$ of items is mapped to the candidate set $C$ of virtual points. We set the weight of every item (a candidate virtual point) to 1 and let $t$ be our target number of virtual points to be added, i.e., $\lambda$. The value of an item, $c_{s}$, is mapped to the loss reduction contributed by the corresponding virtual point. Maximising the values of the items in subset $A$ is mapped to maximising the total loss reduction of the virtual points in $V$. Due to the nature of our problem, the total loss reduction when multiple virtual points are added together varies from the sum of the loss reduction when the virtual points are added individually. This can be represented as:
(3) $|\sum_{s\in A}c_{s}|=r\sum_{s\in A}|c_{s}|,$
where $|c_{s}|$ is the magnitude of the value of an item (a virtual point) and $r\in\mathbb{R}$ is a parameter. Here, $r$ is deterministic since it could be calculated based on Equation 1. As such, the total value of the combined items can be transformed to the sum of the individual item values, and maximising the latter for the Knapsack problem can be mapped to maximising the former for our problem.
The transformation between the two problems can be done in polynomial time since there is a one-to-one mapping between them. Thus, when our problem is solved, the Knapsack problem is also solved in polynomial time. As such, our problem is NP-hard. ∎
@@
Due to the computational complexity in finding an exact optimal solution for the learned index smoothing problem over a large set of keys, next, we consider two practical variants of the problem and propose highly effective heuristic solutions: (1) smoothing the CDF for the subset of keys $K_{i}$ indexed by an indexing function $f_{i}$ (Section 4); and (2) smoothing the CDFs for all $m$ subsets $K_{i}$ when they are indexed under a hierarchical learned index (Section 5).
4. CDF Smoothing for a Single Linear Model